if two angles of the triangle measure 54, what is the measure of the third angle?
Answer by nerdybill(7384)

if two angles of the triangle measure 54, what is the measure of the third angle?
.
The sum of the three interior angles of any triangle is 180 so
.
Let x = measure of third angle
then
x + 54 + 54 = 180
x + 108 = 180
x = 180-108
x = 72 degrees